OpenStudy (chelseafry16):

What is the cube root of -729a^9b^6 a_ -9a^3b^2 b_ -9a^2b^3 c_ -8a^3b^2 d_ -8a^2b^3

OpenStudy (chrishogan):

I can help you. :)

OpenStudy (chelseafry16):

thank you.

OpenStudy (chrishogan):

So first, do you know what a cube root is?

OpenStudy (chelseafry16):

no

OpenStudy (chrishogan):

A cubed root would be solved with \[3\sqrt{x} \] so if x were 343, your cubed root would be 7.

OpenStudy (chrishogan):

For example, the cubed root of 27 is 3 as 3 multiplied by 3 multiplied by 3 is 27. The cubed root of 8,120,601 is 201, as 201 multiplied by 201 multiplied by 201 is 8,120,601. The cubed root of 8 is 2.8284271247461900976033774484194 as 2.8284271247461900976033774484194 multiplied by 2.8284271247461900976033774484194 is 8.

OpenStudy (chrishogan):

Find the cube root of 512 1) Try a number – 5 : 5 times 5 times 5 = 125 (too low) 2) Try another number that is more than 5 – 6 : 6 times 6 times 6 = 216 (too low) 3) Try a number that is more than 6 – 10 – 10 times 10 times 10 = 1,000 (too high) 4) Try a number in between 6 and 10 – 8 – 8 times 8 times 8 = 512 (answer)

OpenStudy (chrishogan):

so what would the answer to this be? \[3\sqrt{729}\]

OpenStudy (chrishogan):

Another way is to use a calculator. and cube 729.

OpenStudy (chrishogan):

Am I going too fast..?

OpenStudy (chelseafry16):

no

OpenStudy (chrishogan):

Alright good. Do you have a calculator on your pc?

OpenStudy (chelseafry16):

yea

OpenStudy (chelseafry16):

i got 81

OpenStudy (chrishogan):

81, isn't quite right.

OpenStudy (chrishogan):

Find view, and click Scientific Calculator.

OpenStudy (chrishogan):

Then put in 729, and find this button. Then, tell me your answer.

OpenStudy (chelseafry16):

its not on my calculator

OpenStudy (chrishogan):

Oh.. Anyways, the answer would be 9, but for our question 729 is negative so we'd get -9. Then you have, \[-9a^9b^6\] so then we divide our exponents by 3, leaving us with \[-9a^3b^2\] which one of your answer choices look like that?

OpenStudy (chelseafry16):

the first one

OpenStudy (chrishogan):

Correct, so that's your answer. :)

OpenStudy (chelseafry16):

thank you!!!!

OpenStudy (chrishogan):

No problem. :)